draw axial force, shear force and bending moment diagrams
2m 4m 30 kN 2m 50 kN 4m A 3 m 1m -3m 1.5m
Answer #1

4 30 VA HA 5-3o-H¢=o (A) - 2oKN Pu - *u 6a-8na - 200 3vc- 4 H -3va-4n, = 6 . +4n -60 320 -3h-8h > 4 0 VCren A) (B) -60 3 66KM O tan(43) = S30 VSFD .o5 C Cria Furce dlagron Teal Ave Contr ve V 2E